OpenStudy (anonymous):

What was the speed of a train that traveled 112.9 miles in 3.7 hours?

OpenStudy (lgbasallote):

\[speed = \frac{distance}{time}\] \[speed = \frac{112.9}{3.7}\] input in the calculator :)

OpenStudy (anonymous):

 Perhaps this triangle would be helpful in the future. Where D= distance S= Speed T= time  So \[distance = speed \times time\] \[Speed = Distance/Time\] \[Time= Distance/Speed\]